OpenStudy (anonymous):

The work of a student to solve a set of equations is shown below: Equation 1: m = 6 + n Equation 2: 2m = 8 + 8n Step 1: –2(m) = –2(6 + n) 2m = 8 + 8n Step 2: –2m = –12 – 2n 2m = 8 + 8n Step 3: –12 – 2n = 8 + 8n Step 4: –20 = 10n Step 5: n = –2 In which step did the student first make an error? (4 points) Step 2 Step 3 Step 4 Step 5

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

i think its step 2

OpenStudy (godlovesme):

Im soo sorry I was afk :/ Just give me one sec

OpenStudy (anonymous):

its ok

OpenStudy (godlovesme):

Im not sure but I think it's 3 :/

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

i have one more i need checked is that ok ??

OpenStudy (godlovesme):

Sure thing :) but I can't stay for another just one more sorry tho :(

OpenStudy (anonymous):

A student is trying to solve the system of two equations given below: Equation P: x + y = 8 Equation Q: 2x + 5y = 24 Which of the following steps can be used to eliminate the x term? (4 points) 2(x + y = 8) –2(x + y = 8) –1(2x + 5y = 24) –2(2x + 5y = 24)

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

i only have this one then im done